OpenStudy (anonymous):

Solve the equation 4x^2-28x+40=0

OpenStudy (anonymous):

Divide through by 4

OpenStudy (gw2011):

4x^2-28x+40=0 4(x^2-7x+10)=0 4(x-5)(x-2)=0 Divide thia equation by 4 and you get: (x-5)(x-2)=0 x-5=0 x=5 x-2=0 x=2 Therefore, x=5 and x=2